OpenStudy (anonymous):

2x - 3y - 4z = -21 -4x + 2y - 3z = -14 -3x - 4y + 2z = -10

OpenStudy (fibonaccichick666):

SO, can you spilt it up into 3 eq for me? HINT: They will all =-10

OpenStudy (anonymous):

2x - 3y - 4z = -21 -4x + 2y - 3z = -14 -3x - 4y + 2z = -10

OpenStudy (fibonaccichick666):

what happened to your -21 in the second?

OpenStudy (fibonaccichick666):

and your -14 in the 3rd?

OpenStudy (fibonaccichick666):

or, is it already 3 eq not one bog long eq?

OpenStudy (anonymous):

it was meant to be on the equation before

OpenStudy (anonymous):

yes its 3 equations

OpenStudy (fibonaccichick666):

ok so there are a few ways to solve a system of three equations

OpenStudy (anonymous):

i tired eliminating and i just cant do it

OpenStudy (anonymous):

tried*

OpenStudy (fibonaccichick666):

1. you solve the others for a variable that is convenient then plug and chug. 2. you turn it into a matrix and put it into reduced row echelon form 3. you add or subtract eq. So, if you know linear algebra2 is the easiest. If you don't 3 is the easiest.

OpenStudy (fibonaccichick666):

Try this for me, add 2 times equation 1 to eq. 2

OpenStudy (anonymous):

what do you mean?

OpenStudy (fibonaccichick666):

ie if I had 4x+2y=6 and -8x+3y=12 I would add 2 times eq 1 to two like so 2(4x+2y)=2(6) 8x+4y=12 so now, + -8x+3y=12 -------------------- 0x+7y=24 so then you get 7y=24 and y=24/3 Do you see?

OpenStudy (anonymous):

its 3 variables though

OpenStudy (fibonaccichick666):

Same idea

OpenStudy (anonymous):

i will try it again then

OpenStudy (fibonaccichick666):

You do the same thing

OpenStudy (fibonaccichick666):

ok

OpenStudy (fibonaccichick666):

Show every step please, even the teeny weeny ones

OpenStudy (anonymous):

alright